OpenStudy (anonymous):

Show work and do not use a calculator: 2^(5/2)-2^(3/2)

OpenStudy (anonymous):

Oops, solved this one for myself 2^(5/2)=2*2^(3/2) 2*2^(3/2)-2^(3/2)=1*2^(3/2)=2^3/2